Grade 11Physical Chemistry

At 300K, 5% glucose and 1.1% Kcl solution are isotonic. Calculate the degree of dissociation of Kcl. Percentage given is in mass/volume

Saurav Raj
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
For Glucose,
1xC1xSxT = ixC2xSxT
C1 = (5/180)*1000/100
and C2 = (1.1/111)*1000/100
 (5/180)*1000/100 = i*(1.1/111)*1000/100
0.099i = 0.277
i = 2.805
 
Degree of dissociation=>
KCl           =============>    K+           +                 Cl-
C0                                             0                                0
C0(1-a)                                     C0a                            C0a
a => (2.805 – 1)*100/2.805 => 64.3%